OpenStudy (anonymous):

graph the equations using the x and y intercepts x/6-y/5=1

jimthompson5910 (jim_thompson5910):

This equation is in intercept form (or close enough to it). Intercept form is x/a + y/b = 1 and the x-intercept is (a, 0) and the y-intercept is (0,b) ------------------------------------------------------- x/6-y/5=1 can be rewritten as x/6+y/(-5)=1 so we can see that a = 6 and b = -5 So the x-intercept is (6,0) and the y-intercept is (0,-5)

jimthompson5910 (jim_thompson5910):

Alternatively, plug in x = 0 and solve for y to get the the y intercept. Plug in y = 0 and solve for x to get the x-intercept. You'll get the same answers.